OpenStudy (anonymous):

If y - (x+3)2, then (-2x6)2 would equal what ?

OpenStudy (anonymous):

You need to know what y - 2(x+3) equals first. Also, what is the second expression? Is that supposed to be 2[(-2)(6)x]?

OpenStudy (anonymous):

The second expression is (-2x-6)2.

OpenStudy (anonymous):

the first expression is (y=(x+3)2

OpenStudy (anonymous):

OK, so that simplifies to -4x - 12 (second expression).

OpenStudy (anonymous):

In solving for x in the first expression, you have to get x to the side that y is on, and vice versa. Thus, y = (x+3)2 would simplify to y = 2x + 6, and then you'd subtract both sides by 6. You get y - 6 = 2x, and further, (y - 6)/2 = x. Since that's what x equals, you can set the x in 2(-2x-6), which simplifies to -4x - 12, to make it -4[(y - 6)/2] - 12. Distribute the -4 so that it's now (-4y+24)/2 - 12. The 2 reduces (-4y+24)/2 - 12 to -2y + 12 - 12. 12 - 12 is 0, so get rid of those. You now have -2y.

OpenStudy (anonymous):

would that be -2ysquared or just -2y ?

OpenStudy (anonymous):

Just -2y.

OpenStudy (anonymous):

I don't have an option for -2y for an answer, just -2y squared.

OpenStudy (anonymous):

Thanks for your help !!!

OpenStudy (anonymous):

No problem!